calculate the gravitational force of attraction between the electron of a hydrogen atom and the position when the electron orbits the proton at its average distance of 0.53×10^-10 take mass of electron as 9×10^-31 kg and mass of proton as 1.67×10^-27kg


1
Expert's answer
2021-03-18T14:03:48-0400
"F=\dfrac{Gm_em_p}{r^2},""F=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot9.1\cdot10^{-31}\ kg\cdot1.67\cdot10^{-27}\ kg}{(0.53\cdot10^{-10}\ m)^2},""F=3.61\cdot10^{-47}\ N."